A veternarian needs a mixture that contains 12% of a certain mdeication to treat an injured bird. How many milliliters of a 16% solution should be mixed with 6mL of a 7% solution to obtain a solution that is 12% medication?
Answer by checkley77(12844)

.16X+.07*6=.12(X+6)
.16X+.42=.12X+.72
.16X-12X=.72-.42
.04X=.30
X=.30/.04
X=7.5 ML OF 16% SOLUTION IS USED.
PROOF:
.16*7.5+.07*6=.12(7.5+6)
1.2+.42=.12*13.5
1.62=1.62